At given place acceln due to gravity is g m/s* a sphere of lead of density d is gently released in column of liquid of density €. If d> €The sphere will fall with acceleration

Grade:12

1 Answers

DR STRANGE
7 years ago
Hi Nizam,
buoyant force (upward) = V*€*g     V=volume of sphere
force due to gravity (downward) =Mg= d*V*g …........... (   M=V*d )
net force F=Ma=d*V*a = downward – upward=V*g*(d-€)
hence a=\frac{g(d-\euro )}{d}=g(1-\frac{\euro }{d})
Plz approve if you are satisfied :)